OpenStudy (4meisu):

Consider the infinite geometric sequence 3000,-1800,1080,-648,.. Find the common ratio, the 10th term, and find the exact sum of the infinite sequence.

OpenStudy (anonymous):

is the common ratio -3/5

OpenStudy (anonymous):

10th term = \[3000*(-3/5)^{9}\]

OpenStudy (callisto):

common ratio = 2nd term / first term = -1800/3000 = -3/5 Sum of GS to infinity = a/(1-r) = 3000 / [1-(-3/5)] = can you work it out?